# Renormalization group approach to graphene bilayers

### v1

@@ III.3.3 Physical discussion: origin of the combination $\ell^{2}(\lambda+2\mu)/2$
We now discuss the origin of the combination $\ell^{2}(\lambda+2\mu)/2$ that appears in the effective bending rigidity $\kappa_{\mathrm{eff}}$. This contribution originates from the interlayer shear coupling, which converts in-plane compression and shear of the two coupled layers into an additional bending energy for the mean membrane. We now show this in more detail.
We first recall the physical situation. The interlayer distance is strictly fixed by the large-$g_{1}$ constraint $\boldsymbol{s}^{2}\simeq\ell^{2}$. The term proportional to $g_{2}$ also plays a crucial geometric role. Writing the relative field as
$\boldsymbol{s}=s_{n}\,\boldsymbol{n}+\boldsymbol{s}_{\parallel},$ (87)
with $\boldsymbol{n}$ the local normal to the mean surface and $\boldsymbol{s}_{\parallel}$ its tangential component, one has, for small slopes, $\boldsymbol{s}\cdot\partial_{\alpha}\boldsymbol{r}\simeq\boldsymbol{s}_{\parallel}\cdot\partial_{\alpha}\boldsymbol{r}$. The contribution
$\frac{g_{2}}{2\ell^{2}}\,\big(\boldsymbol{s}\cdot\partial_{\alpha}\boldsymbol{r}\big)^{2}$ (88)
therefore penalizes precisely the tangential component $\boldsymbol{s}_{\parallel}$, i.e. lateral sliding between the two layers at fixed normal separation. In the regime where $g_{2}$ is large, the relative field is thus effectively constrained to be normal to the mean surface, $\boldsymbol{s}\simeq\ell\,\boldsymbol{n}$, so that the bilayer behaves as two parallel membranes, separated by a fixed distance $\ell$, without in-plane slip.
We now consider a purely bending deformation of the bilayer and make this explicit by working in the Monge representation. We parameterize the mean surface as
$\boldsymbol{r}(x,y)=(x,\,y,\,h(x,y)),$ (89)
and, for simplicity, we first restrict to a uniaxial bending mode with $h=h(x)$ and small slopes, $|\partial_{x}h|\ll 1$. In this setting, the mean surface acts as a neutral surface for bending, while the two physical layers 1 and 2, parametrized by the vectors
$\boldsymbol{r}_{1}=\boldsymbol{r}+\frac{\ell}{2}\,\boldsymbol{n},\qquad\boldsymbol{r}_{2}=\boldsymbol{r}-\frac{\ell}{2}\,\boldsymbol{n}$ (90)
experience opposite in-plane strains along the bending direction.
To make this more precise, we use the expression of the metric $g_{\alpha\beta}(d)$ induced at a distance $d$ from a reference surface with metric $g_{\alpha\beta}$. One has, see App. D
$g_{\alpha\beta}(d)=g_{\alpha\beta}-2d\,K_{\alpha\beta}+d^{2}(K^{2})_{\alpha\beta},$ (91)
@@
where $K_{\alpha\beta}$ is the second fundamental form and $(K^{2})_{\alpha\beta}$ the corresponding quadratic curvature tensor.
In our case one has $d=+\ell/2$ and $d=-\ell/2$. Plugging these values into Eq. (91) and expanding to first order in $\ell$, we obtain
$g^{(1,2)}_{\alpha\beta}=g_{\alpha\beta}\mp\ell\,K_{\alpha\beta}+O(\ell^{2}),$ (92)
where the sign corresponds to the choice $\pm\ell/2$ along the normal $\boldsymbol{n}$ and where the indices and 2 refer to the corresponding surfaces.
The strain tensor on each surface is defined (with our conventions) as
$u_{\alpha\beta}^{(i)}=\frac{1}{2}\big(g_{\alpha\beta}^{(i)}-\delta_{\alpha\beta}\big),\qquad i=1,2.$ (93)
Let $g_{\alpha\beta}^{(0)}:=g_{\alpha\beta}$ and
$u_{\alpha\beta}^{(0)}=\frac{1}{2}\big(g_{\alpha\beta}^{(0)}-\delta_{\alpha\beta}\big)$ (94)
be the strain of the mean surface. Using Eq. (92), we can write
$u_{\alpha\beta}^{(1,2)}=u_{\alpha\beta}^{(0)}\mp\frac{\ell}{2}\,K_{\alpha\beta}+O(\ell^{2})\,.$ (95)
For a pure bending mode of the mean surface (neutral surface), we can neglect $u_{\alpha\beta}^{(0)}$ and keep only the curvature-induced part:
$u_{\alpha\beta}^{(1)}\simeq-\frac{\ell}{2}\,K_{\alpha\beta},\qquad u_{\alpha\beta}^{(2)}\simeq+\frac{\ell}{2}\,K_{\alpha\beta}\ .$ (96)
The in-plane elastic energy density of a single layer with Lamé coefficients $\lambda,\mu$ is
$f_{\text{el}}=\lambda\,u_{\alpha\alpha}^{2}+2\mu\,u_{\alpha\beta}^{2}\ .$ (97)
For the uniaxial bending mode described above (say $h=h(x)$ in Monge gauge), one principal curvature direction dominates, and we can effectively take
$u_{xx}^{(1,2)}\simeq\pm\frac{\ell}{2}\,\kappa,\qquad u_{yy}^{(1,2)}\simeq u_{xy}^{(1,2)}\simeq 0,$ (98)
where $\kappa\sim\partial_{x}^{2}h$ is the local curvature.
In this uniaxial case,
$u_{\alpha\alpha}=u_{xx}=\varepsilon,\qquad u_{\alpha\beta}^{2}=u_{xx}^{2}=\varepsilon^{2},$ (99)
so the elastic energy for each sheet is
$f_{\text{el}}^{(i)}=(\lambda+2\mu)\,(\varepsilon^{(i)})^{2}.$ (100)
Substituting $\varepsilon^{(1,2)}=\pm(\ell/2)\,\kappa$, we obtain
$f_{\text{el}}^{(1)}=(\lambda+2\mu)\left(\frac{\ell}{2}\kappa\right)^{2},\qquad f_{\text{el}}^{(2)}=(\lambda+2\mu)\left(-\frac{\ell}{2}\kappa\right)^{2},$ (101)
and the total elastic energy density associated with bending of the bilayer is
$f_{\text{el}}^{\text{(bilayer)}}$ $=f_{\text{el}}^{(1)}+f_{\text{el}}^{(2)}$
$=2(\lambda+2\mu)\,\frac{\ell^{2}}{4}\,\kappa^{2}$
$=\frac{\ell^{2}}{2}(\lambda+2\mu)\,\kappa^{2}.$ (102)
In Monge gauge, $\kappa\sim\partial_{x}^{2}h$, so this is a contribution of the form
$\frac{\ell^{2}}{2}(\lambda+2\mu)\,(\partial^{2}h)^{2}$ (103)
in the effective average action for the mean surface. It is thus natural to interpret
$\kappa_{\text{el}}=\frac{\ell^{2}}{2}(\lambda+2\mu)$ (104)
as the elastic bending rigidity induced by the in-plane elasticity of the two layers, once the bilayer geometry $\boldsymbol{r}\pm\ell\boldsymbol{n}/2$ is enforced by the $g_{1}, large$-$g_{2}$ limit.
@@ III.3.4 RG equations in the flat phase
We are only interested in the flat phase of the system. We can therefore take, prior to the calculation, the limit $\bar{\zeta}_{k}\gg 1$. The RG equations that control the infrared behaviour of membrane bilayers are then given by:
@@
$P_{c_{k}}(\boldsymbol{q})=(Z_{k}+c_{k})\,\boldsymbol{q}^{4}+R_{k}(\boldsymbol{q})\,.$ (112)
@@ IV Physical results
We now turn to a discussion of our results. We begin in Sec. IV A with a qualitative analysis of the short- and long-distance behaviour of the effective bending rigidity, from which we obtain an estimate of the mechanical crossover scale $k_{c}$. In Sec. IV B, we specify the initial conditions of the RG flow for realistic (graphene) parameter values. We then discuss, in Sec. IV C, the Ginzburg scale $k_{G}$ associated with the harmonic-anharmonic crossover. Finally, in Sec. IV D, we study the flow of the running anomalous dimension and Young modulus.
@@ IV.1 Mechanical crossover scale
Here we analyze the short- and long-distance behaviour of the effective bending rigidity and provide an estimate of the mechanical crossover length scale.
@@
It is worth stressing at this point that our NPRG approach naturally provides a scale-dependent effective bending rigidity $\kappa_{\text{eff}}(k)$, while most experimental determinations of $\kappa$ extract an effective value averaged over a finite wave-vector window set by the sample geometry and the probing technique (nanoindentation, snap-through, modal analysis, or scattering). In this sense, measurements typically probe $\kappa_{\text{eff}}(k\simeq k_{\text{geom}})$, with $k_{\text{geom}}\sim 1/L$ the inverse size of the probed membrane or the typical wave vector of the relevant mode, rather than the strict $k\to 0$ limit. The RG description in terms of $\kappa_{\text{eff}}(k)$ is therefore particularly well suited to interpret these experiments, since it directly encodes how the apparent bending rigidity changes when the probed length scale is varied.
@@
IV.1.1 Short- and long-distance behaviour of the effective bending rigidity
@@ IV.1.2 Mechanical crossover scale : first analysis
Under this approximation, Eq. (129) gives the simple estimate
@@
Numerically, with the bare values of graphene [45] $\kappa_{*}=\kappa_{g}=1,$ and $\ell=\ell_{g}=3.25\,\text{A}$ and the fixed point values [40]: $\bar{\mu}_{*}\simeq 6.21,\bar{\lambda}_{*}\simeq-3.10$, $\eta=0.85$ and [40] one gets: $k_{c}\simeq 0.06\text{A}^{-1}$ and the RG-time $t_{c}=-\ln k_{c}a=1.88$.
@@
$k_{c}\sim a^{-1}\left[\frac{4\,\bar{\kappa}}{\ell^{2}\big(\bar{\lambda}+2\bar{\mu}\big)}\right]^{\!\frac{1}{2-\eta}}.$ (134)
@@
one gets: $k_{c}\simeq 0.012\text{A}^{-1}$ and the RG-time $t_{c}=-\ln k_{c}a=3.54$.
@@ IV.1.4 Mechanical crossover scale : RG analysis
We finally provide an estimate of the mechanical crossover scale $k_{c}$ based on an RG analysis. We define the crossover RG-time $t_{c}$ as the scale at which $\bar{c}_{k_{c}}\simeq 1$, see Eq.(130). For a temperature $T=500$ K, one finds $k_{c}\simeq 0.0028\,\text{A}^{-1}$ and $t_{c}\simeq 4.96$, see Fig. 3, which is larger than both $t_{1c}$ and $t_{2c}$.
As seen in Fig. 3, the typical crossover RG-time $t_{c}$ lies well above the two scales $t_{1c}$ and $t_{2c}$, see Eq.(138), between which the rigidity crossover would naively be expected to occur. The origin of this discrepancy is the following. When $g_{2}$ is kept finite, the mechanical crossover observed in Ref. [45] is driven by interlayer shear, characterized by $g_{2}$, and naturally takes place between $t_{1c}$ and $t_{2c}$. In the present case, where we consider the limit of large $g_{2}$, we define the crossover scale by
@@
Figure 3: Behaviour of $\bar{c}_{k}$ – dashed-dot curve – as a function of $t=-\ln ka$ for $T=500$ K. One has displayed the typical RG-time $t_{c}$ (dashed-dot vertical line) associated with the bending rigidity crossover and the RG-times $t_{1c}$ and $t_{2c}$ (full vertical lines).
@@ IV.2 Initial conditions RG flow: graphene parameters
which is much larger than unity. We are therefore precisely in the conditions required to observe a mechanical crossover.
@@ IV.3 Ginzburg scale
At a temperature T=500T=500K one finds $k_{G}=1.82$, as defined in Eq.(150), and $t_{G}=1.50$, that lies between $t_{1G}$ and $t_{2G}$ as defined in Eq.(146), see Fig.(4). This turns out to be true at all temperatures. From now on, we will consider $t_{G}$ as the harmonic-to-anharmonic crossover RG-time scale.
Figure 4: Behaviour of $\eta_{k}$ – dashed curve – as a function of $t=-\ln ka$ for $T=500$ K; One gives also the typical Ginzburg RG-time scale $t_{G}$ (dashed vertical line) and the bare Ginzburg RG-time scales $t_{G1}$ and $t_{G2}$ (full vertical lines).
@@ IV.4 RG analysis
We now present our results, which concern the RG evolution of both the running anomalous dimension $\eta_{k}$, given by Eq. (116) and the running Young modulus $\bar{Y}_{k}=4\bar{\mu}_{k}(\bar{\lambda}_{k}+\bar{\mu}_{k})/(\bar{\lambda}_{k}+2\bar{\mu}_{k})$, given by Eq. (115), for various temperatures. We have considered $T=10\,\text{K}$, $T=100\,\text{K}$ and $T=1000\,\text{K}$. In each case, we indicate the Ginzburg scale $t_{G}$ by a vertical dotted line and the mechanical crossover scale $t_{c}$ by a vertical dot-dashed line.
@@
IV.4.1 Behaviour of the anomalous dimension $\eta_{k}^{B}$
In Fig.(5), (6) and (7) we have represented the anomalous dimensions of a bilayer $\eta_{k}^{B}$ and that of a monolayer $\eta_{k}^{m}$ at the temperatures $T=10\text{K}$, $T=100\text{K}$ and $T=1000\text{K}$. First, one observes that at large times, the RG flow of the anomalous dimension of a bilayer converges towards that of a monolayer that, itself, reaches the value taken at the IR fixed point of the flat phase: $\eta\simeq 0.85$. Second, one sees that harmonic-to-anharmonic crossover takes place long before the mechanical one, and that, at all temperatures. Third, one observes that the two crossovers occur at earlier and earlier RG-times when the temperature increases, as expected.

### v2

@@ III.3.3 Physical discussion: origin of the combination $\ell^{2}(\lambda+2\mu)/2$
We now discuss the origin of the combination $\ell^{2}(\lambda+2\mu)/2$ appearing in the effective bending rigidity $\kappa_{\mathrm{eff}}$. It reflects the additional in-plane elastic cost generated by bending when the two sheets are constrained to stay at fixed separation $\ell$ and to bend without lateral slip. The large-$g_{1}$ constraint enforces $|\boldsymbol{s}|\simeq\ell$, while the $g_{2}$ term penalizes the tangential component $\boldsymbol{s}_{\parallel}$; thus, for large $g_{2}$, $\boldsymbol{s}\simeq\ell\,\boldsymbol{n}$ and the two physical layers are parallel surfaces,
$\boldsymbol{r}_{1,2}=\boldsymbol{r}\pm\frac{\ell}{2}\boldsymbol{n}.$ (85)
where the indices 1 and 2 refer to bilayer 1 and 2.
Using the parallel-surface metric expansion (App. E)
$g_{\alpha\beta}(d)=g_{\alpha\beta}-2d\,K_{\alpha\beta}+{\cal O}(d^{2}),$ (86)
@@
where $K_{\alpha\beta}$ is the second fundamental form. In our case one has $d=+{\ell/2}$ and $d=-{\ell/2}$ and the layer metrics at $d=\pm\ell/2$ read
$g^{(1,2)}_{\alpha\beta}=g_{\alpha\beta}\mp\ell\,K_{\alpha\beta}+{\cal O}(\ell^{2}),$ (87)
hence the corresponding strains
$u^{(1,2)}_{\alpha\beta}=\frac{1}{2}\big(g^{(1,2)}_{\alpha\beta}-\delta_{\alpha\beta}\big)=u_{\alpha\beta}(0)\mp\frac{\ell}{2}K_{\alpha\beta}+{\cal O}(\ell^{2}),$ (88)
where $u_{\alpha\beta}(0)\equiv\tfrac{1}{2}(g_{\alpha\beta}-\delta_{\alpha\beta})$ is the mean-surface strain.
The in-plane elastic energy density of one sheet is
$f_{\rm el}=\lambda\,(u_{\gamma\gamma})^{2}+2\mu\,u_{\alpha\beta}u_{\alpha\beta}.$ (89)
Substituting (88) and summing the two layers, all terms linear in $K_{\alpha\beta}$ cancel by symmetry, and one obtains
$\begin{array}[]{ll}f_{\rm el}^{\rm(bilayer)}&=\displaystyle\sum_{i=1,2}\Big[\lambda\,(u^{(i)}_{\gamma\gamma})^{2}+2\mu\,u^{(i)}_{\alpha\beta}u^{(i)}_{\alpha\beta}\Big]\\ \\ &=2\Big[\lambda\,u_{\gamma\gamma}(0)^{2}+2\mu\,u_{\alpha\beta}(0)u_{\alpha\beta}(0)\Big]\\ \\ &+\displaystyle\frac{\ell^{2}}{2}\Big[\lambda\,(K_{\gamma\gamma})^{2}+2\mu\,K_{\alpha\beta}K_{\alpha\beta}\Big]+{\cal O}(\ell^{4}).\end{array}$ (90)
The first bracket is simply the in-plane elastic energy of the mean surface. The second bracket is the curvature-induced contribution: it is quadratic in $K_{\alpha\beta}$ and therefore renormalizes the bending sector of the effective theory.
For a uniaxial bending mode (Monge gauge, $h=h(x)$), the dominant component is $K_{xx}\simeq\kappa$, with $K_{yy}\simeq K_{xy}\simeq 0$, so that
$(K_{\gamma\gamma})^{2}=K_{xx}^{2}=\kappa^{2},\qquad K_{\alpha\beta}K_{\alpha\beta}=K_{xx}^{2}=\kappa^{2},$ (91)
and the curvature part of (90) reduces to
$f_{\rm el,curv}^{\rm(bilayer)}=\frac{\ell^{2}}{2}(\lambda+2\mu)\,\kappa^{2}\;\sim\;\frac{\ell^{2}}{2}(\lambda+2\mu)\,(\partial_{x}^{2}h)^{2}.$ (92)
This identifies the induced elastic bending rigidity
$\kappa_{\rm el}=\frac{\ell^{2}}{2}(\lambda+2\mu)$ (93)
which contributes additively to $\kappa_{\mathrm{eff}}$.
@@ III.3.4 RG equations in the flat phase
We are only interested in the flat phase of the system. We can therefore take, prior to the calculation, the limit $\bar{\zeta}_{k}\gg 1$. The RG equations that control the infrared behavior of membrane bilayers are then given by:
@@
$P_{c_{k}}(\boldsymbol{q})=(Z_{k}+c_{k})\,\boldsymbol{q}^{4}+{\cal R}_{k}(\boldsymbol{q})\,.$ (101)
@@ IV Physical results
We now turn to a discussion of our results. We begin in Sec. IV A with a qualitative analysis of the short- and long-distance behavior of the effective bending rigidity, from which we obtain an estimate of the mechanical crossover scale $k_{c}$. In Sec. IV B, we specify the initial conditions of the RG flow for realistic (graphene) parameter values. We then discuss, in Sec. IV C, the Ginzburg scale $k_{G}$ associated with the harmonic-anharmonic crossover. Finally, in Sec. IV D, we study the flow of the running anomalous dimension and Young modulus.
@@ IV.1 Mechanical crossover scale
Here we analyze the short- and long-distance behavior of the effective bending rigidity and provide an estimate of the mechanical crossover length scale.
@@
It is worth stressing at this point that our NPRG approach naturally provides a scale-dependent effective bending rigidity $\kappa_{\text{eff}}(k)$, while most experimental determinations of $\kappa$ extract an effective value averaged over a finite wave-vector window set by the sample geometry and the probing technique (nanoindentation, snap-through, modal analysis, or scattering). In this sense, measurements typically probe $\kappa_{\text{eff}}(k\simeq k_{\text{geom}})$, with $k_{\text{geom}}\sim 1/L$ the inverse size of the probed membrane or the typical wave vector of the relevant mode, rather than the strict $k\to 0$ limit. More precisely, in a finite sample of lateral size $L$, the infrared sector is cut off by the smallest accessible wave vector $q_{\min}\sim\alpha/L$, where the numerical factor $\alpha$ depends on the geometry and boundary conditions (e.g. $\alpha=2\pi$ for periodic boundaries, while for a circular drum $\alpha$ is set by the relevant Bessel-mode wave number). Therefore the identification $k\sim 1/L$ is meant at the scaling level and the proportionality constant is not universal; it can be fixed for a given experimental (or numerical) setup either by matching to the dominant eigenmode(s) of the corresponding geometry, or empirically by calibrating $\alpha$ from measurements/simulations where the probed wave-vector window is known. The RG description in terms of $\kappa_{\text{eff}}(k)$ is therefore particularly well suited to interpret these experiments, since it directly encodes how the apparent bending rigidity changes when the probed length scale is varied.
@@
IV.1.1 Short- and long-distance behavior of the effective bending rigidity
@@ IV.1.2 Mechanical crossover scale : first analysis
Under this approximation, Eq. (118) gives, by restoring the dimension of $\ell$, the simple estimate
@@
[…]
@@
$k_{c}\sim a^{-1}\left[\frac{4\,\kappa}{\ell^{2}\big(\lambda+2\mu\big)}\right]^{\!\frac{1}{2-\eta}}.$ (123)
@@
Note that, when the crossover scale $k_{c}$ is estimated from the fixed-point values of the dimensionless couplings, Eq.(122) the microscopic lattice spacing $a$ does not appear explicitly: the condition defining $k_{c}$ compares two contributions to $\kappa_{\mathrm{eff}}(k)$ at the same running scale $k$ and is therefore formulated entirely in terms of renormalized, dimensionless quantities. The dependence on the ultraviolet cutoff $\Lambda\sim a^{-1}$ enters only through the choice of origin of the RG “time” $t=\ln(k/\Lambda)$ and thus drops out of the crossover condition itself, in contrast with bare estimates, Eq.(123), written directly in terms of microscopic parameters.
@@ IV.1.4 Mechanical crossover scale : RG analysis
[…]
As seen in Fig. 3, the typical crossover RG-time $t_{c}$ lies well above the two scales $t_{1c}$ and $t_{2c}$, see Eq.(127), between which the rigidity crossover would naively be expected to occur. The origin of this discrepancy is the following. When $g_{2}$ is kept finite, the mechanical crossover observed in Ref. [45] is driven by interlayer shear, characterized by $g_{2}$, and naturally takes place between $t_{1c}$ and $t_{2c}$. In the present case, we define the crossover scale by
@@
Figure 3: Behavior of $\bar{c}_{k}$ – dashed-dot curve – as a function of $t=-\ln ka$ for $T=500$ K. One has displayed the typical RG-time $t_{c}$ (dashed-dot vertical line) associated with the bending rigidity crossover and the RG-times $t_{1c}$ and $t_{2c}$ (full vertical lines).
@@ IV.2 Initial conditions RG flow: graphene parameters
Let us remark that the present dimensionless formulation, the RG flow equations are temperature independent, see e.g. Eq.(97) while the temperature $T$ enters through the microscopic initial conditions at $k=\Lambda$, since restoring physical units amounts to setting the initial dimensionless couplings proportional to $\beta=1/k_{B}T$.
@@
see Fig.3, which is much larger than unity. We are therefore precisely in the conditions required to observe a mechanical crossover.
@@ IV.3 Ginzburg scale
[…]
Figure 4: Behavior of $\eta_{k}$ – dashed curve – as a function of $t=-\ln ka$ for $T=500$ K; One gives also the typical Ginzburg RG-time scale $t_{G}$ (dashed vertical line) and the bare Ginzburg RG-time scales $t_{1G}$ and $t_{2G}$ (full vertical lines).
@@ IV.4 RG analysis
We now present our results, which concern the RG evolution of both the running anomalous dimension $\eta_{k}$, given by Eq. (105) and the running Young modulus $\bar{Y}_{k}$, given by Eq. (104), for various temperatures. We have considered $T=10\,\text{K}$, $T=100\,\text{K}$ and $T=1000\,\text{K}$. In each case, we indicate the Ginzburg scale $t_{G}$ by a vertical dotted line and the mechanical crossover scale $t_{c}$ by a vertical dot-dashed line.
@@
IV.4.1 Behavior of the anomalous dimension $\eta_{k}^{B}$
In Fig.5-7 we have represented the anomalous dimensions of a bilayer $\eta_{k}^{B}$ and that of a monolayer $\eta_{k}^{m}$ at the temperatures $T=10\text{K}$, $T=100\text{K}$ and $T=1000\text{K}$. First, one observes that at large times, the RG flow of the anomalous dimension of a bilayer converges towards that of a monolayer that, itself, reaches the value taken at the IR fixed point of the flat phase: $\eta\simeq 0.85$. Second, one sees that harmonic-to-anharmonic crossover takes place long before the mechanical one, and that, at all temperatures. Third, one observes that the two crossovers occur at earlier and earlier RG-times when the temperature increases, as expected.